# Representation biases: will we achieve complete understanding by analyzing representations?

### v1

@@ 1.1 Background
In a foundational paper, Marr (1982) famously posited three levels of analysis for understanding an information processing system—computational, algorithmic, and implementational. In the years since, there has been substantial discussion about the (non)independence of the levels, their interactions and joint constraints, and what they imply about representation (e.g. McClamrock, 1991; Churchland and Sejnowski, 1990; O’Brien and Opie, 2009; Stevens, 2012; Piccinini and Shagrir, 2014; Hardcastle and Hardcastle, 2015). In the philosophical literature, similarly, the question of the relationship between representation and computation has been somewhat fraught (e.g. Eliasmith and Anderson, 2003; Piccinini, 2008; Shea, 2018).
@@ 3 How might feature biases impact neuroscience?
For example, a growing area of research relies on comparing representations between AI models and humans or animals (Sucholutsky et al., 2023), and there have been explicit calls for studying representational as well as behavioral alignment of models to the brain (e.g. Feather et al., 2025)—as well as collaborative benchmarks for this purpose (e.g Schrimpf et al., 2020). However, if the AI models are biased towards representing certain features more strongly, this could lead to biased inferences from these analyses and benchmarks. For example, such biases might help to explain why there is surprisingly little difference between different AI model architectures and training objectives in how closely they align with brain representations (Conwell et al., 2023)—if the representations of all the AI models tend to be dominated by the simplest (or most prevalent) features which any objective will learn.
@@
4 Homomorphic encryption: strongly dissociating representation & computation
In the experiments described above, the role that the representations played in the computations of the system was relatively straightforward, even where the representations were biased towards one feature. However, this does not have to be the case. We illustrate this with a final conceptual case study of the possibility for strong dissociation between representation and computation: homomorphic encryption (Gentry, 2009; Van Dijk et al., 2010). While the field of cryptography is largely focused on creating representations that preserve information yet are not easily decodable, in homomorphic encryption schemes it is additionally possible to perform arbitrary computations (any algebraic circuit) over this information while it is encrypted. That is, at each step of such a computation, a new encrypted representation is produced that corresponds to the result of encrypting the representation at that step of the original computation.
This example shows that it is not necessary for a computational system to have any straightforward (e.g. linearly decodable) representation of the features that it uses in its computations. Systematic computations can be performed even over representations that are deliberately crafted to thwart attempts to understand (decrypt) their content.
As a special case, this example also illustrates that systematic compositional computations are possible without requiring representations that are straightforwardly compositional. Encrypted representations are compositional only in the sense that “with the right highly-nonlinear decoding scheme compositional representations can be extracted”—which is also true of some coding schemes typically interpreted as non-compositional, such as idiosyncratic representations of each input. This raises questions about if and when it is feasible to rigorously confirm whether a system’s computations are compositional from representational analyses.
This example illustrates how opaque the relationship between representation and computation can be in the worst case. However, computing using homomorphic encryption is generally too slow for many practical uses, which perhaps hints at one of the constraints that pushes neural and machine learning systems toward simpler representation formats; we will return to this question below.
@@ Acknowledgments
We thank Greta Tuckute and Nathaniel Daw for helpful comments and suggestions on earlier drafts of this paper. We also thank Talia Konkle & the Konklab, Daniel Yamins & the NeuroAILab, and Nicolaus Kriegeskorte & the Zuckerman Mind Brain Behavior Institute for stimulating discussions of these issues. We thank Alex Ku and Murray Shanahan for support.

### v2

@@ 1.1 Background
In a foundational paper, Marr (1982) famously posited three levels of analysis for understanding an information processing system—computational, algorithmic, and implementational. In the years since, there has been substantial discussion about the (non)independence of the levels, their interactions and joint constraints, and what they imply about representation (e.g. McClamrock, 1991; Churchland and Sejnowski, 1990; O’Brien and Opie, 2009; Stevens, 2012; Piccinini and Shagrir, 2014; Hardcastle and Hardcastle, 2015). In the philosophical literature, similarly, the question of the relationship between representation and computation has been somewhat fraught (e.g. Eliasmith and Anderson, 2003; Piccinini, 2008; Van Gelder, 2014; Shea, 2018).
@@ 3 How might feature biases impact neuroscience?
For example, a growing area of research relies on comparing representations between AI models and humans or animals (Sucholutsky et al., 2023), and there have been explicit calls for studying representational as well as behavioral alignment of models to the brain (e.g. Feather et al., 2025)—as well as collaborative benchmarks for this purpose (e.g Schrimpf et al., 2020). However, if the AI models are biased towards representing certain features more strongly, this could lead to biased inferences from these analyses and benchmarks. For example, such biases might help to explain why there is surprisingly little difference between different AI model architectures in how closely they align with brain representations (Conwell et al., 2023)—if the representations of the various models tend to be dominated by the simplest (or most prevalent) features which any objective will learn.
@@
4 Homomorphic encryption: strongly dissociating computation from superficial patterns of representation
In the experiments described above, the role that the representations played in the computations of the system was relatively straightforward, even where the representations were biased towards one feature. However, this does not have to be the case. We illustrate this with a final conceptual case study of the possibility for strong dissociation between computation and patterns of representation: homomorphic encryption (Gentry, 2009; Van Dijk et al., 2010). While the field of cryptography is largely focused on creating representations that preserve information yet are not easily decodable, in homomorphic encryption schemes it is additionally possible to perform arbitrary computations (any algebraic circuit) over this information while it is encrypted. That is, at each step of such a computation, a new encrypted representation is produced that, when decrypted, yields the representation at that step of the original computation.
This example shows that it is not necessary for a computational system to have any straightforward (e.g. linearly decodable) format of representation of the features that it uses in its computations. Systematic computations can be performed even over representations that are deliberately crafted to thwart attempts to understand (decrypt) their content.
As a special case, this example also illustrates that systematic compositional computations are possible without requiring representations that are straightforwardly (e.g. contatenatively) compositional in their original format. Encrypted representations are compositional only in the sense that “with the right highly-nonlinear decoding scheme compositional representations can be extracted”—which is also true of some coding schemes typically interpreted as non-compositional, such as idiosyncratic representations of each input. That is, this case provides a clean example of a functionally-but-not-concatenatively compositional representation (cf. Van Gelder, 2014). This raises questions about if and when it is feasible to rigorously confirm whether a system’s computations are compositional from representational analyses.
This example illustrates how opaque the relationship between the superficial format of representation and computation can be in the worst case. However, computing using homomorphic encryption is generally too slow for many practical uses, which perhaps hints at one of the practical constraints that pushes neural and machine learning systems toward simpler—and more aligned—representation formats; we will return to this question below.
@@ Acknowledgments
We thank Greta Tuckute, Nathaniel Daw, Lukas Muttenthaler, and Raphaël Millière for helpful comments and suggestions on earlier drafts of this paper. We also thank Talia Konkle & the Konklab, Daniel Yamins & the NeuroAILab, and Nicolaus Kriegeskorte & the Zuckerman Mind Brain Behavior Institute for stimulating discussions of these issues. We thank Alex Ku and Murray Shanahan for support.